OpenStudy (anonymous):

Consider the function y = 3x5 – 25x3 + 60x + 1. Use the first or second derivative test to test the critical points. How many relative maxima did you find?

OpenStudy (campbell_st):

so what did you get for f'(x) and f"(x)

OpenStudy (anonymous):

15x4-75x2+60=f' 60x3-150x=f"

OpenStudy (anonymous):

i think i found the critical points already. but im confused on how to test it to find the maxima.

OpenStudy (anonymous):

critical points: –2, –1, 1, 2...??? i think.

OpenStudy (campbell_st):

ye, thats correct, so now test them in the 2nd derivative by substituting if f"(a) > 0 you have a minimum f"(a) = 0 a horizontal point of inflexion f"(a) < 0 you have a maximum... so text the solutions and you should be able to find the maxima... hope it helps

OpenStudy (anonymous):

so tests the critical pts in the f'' and follow those rules...?

OpenStudy (campbell_st):

oops a is a solution to the 1st derivative

OpenStudy (campbell_st):

yes... for example f"(-1) = 60(-1)^3 - 150(-1) = 90 so a minimum...

OpenStudy (anonymous):

that doesnt make sense. sorry, youre good at guiding me, but im having trouble embedding what youre saying with what ive learned. doest a min have to have a negative numbers?

OpenStudy (anonymous):

YES!!! i solved it! it has 2 maximum!!! right?

OpenStudy (campbell_st):

no a minimum can be any where on the graph because there are relative to each other.. here is a graph of your curve

OpenStudy (campbell_st):

thats correct... at x = -2 and x = -1

OpenStudy (anonymous):

does the original equation have a min at f' when x=1

OpenStudy (anonymous):

i mean max

OpenStudy (campbell_st):

as you can see from the graphs it has mins at x = -1 and x = 2 you can solve the 2nd derivative for possible points of inflexion, and check by testing either side of the points for a change in concavity

OpenStudy (campbell_st):

oopps yes... a max at x = 1.... not -1

OpenStudy (anonymous):

whats the difference between a local and relative min/max

OpenStudy (anonymous):

nevermind! i just checked my answers and i got them all right! thank you!